OpenStudy (anonymous):

A wooden pyramid, 9 inches tall, has a square base. A carpenter increases the dimensions of the wooden pyramid by a factor of 4 and makes a larger pyramid with the new dimensions. Describe in complete sentences the ratio of the volumes of the two pyramids.

Directrix (directrix):

The two pyramids are similar with scale factor 1/4. If two polyhedra are similar then the cube of their scale factor is the ratio of their volumes. (1/4)^3 = V/ V' where V is the volume of the original pyramid and V' is the volume of the enlarged pyramid. 1/64 = V/V' The volume of the new pyramid is 64 times that of the original one.

OpenStudy (anonymous):

thanks